if a number is added to its square, the result is 56. Find the number.
Answer by user_dude2008(1862)

x+x^2=56

x^2+x-56=0

(x+8)(x-7)=0

x+8=0 or x-7=0

x=-8 or x=7

Numbers are 8 or -7